# Jailbreaking Prompt Attack: A Controllable Adversarial Attack against Diffusion Models

### v3

@@ Abstract
We perform extensive experiments with open-sourced T2I models, e.g. (CompVis 2022a) and closed-sourced online services, e.g. (Ramesh et al. 2022; MidJourney 2023; Podell et al. 2023) with black-box safety checkers. Results show that (1) JPA bypasses both text and image safety checkers (2) while preserving high semantic alignment with the target prompt. (3) JPA demonstrates a much faster speed than previous methods and can be executed in a fully automated manner. These merits render it a valuable tool for robustness evaluation in future text-to-image generation research.
@@
Introduction
@@
To mitigate these risks, contemporary T2I models incorporate stringent safety checkers and removal techniques to prevent the generation of NSFW content, including text-based (Jieli 2023; George 2023), image-based (Laion-ai 2023; Chhabra 2020) and (Rando et al. 2022; CompVis 2022b) text-image-based safety checkers. These checkers either filter out sensitive words or classify unsafe images. Concept removal methods (Schramowski et al. 2023; Gandikota et al. 2023; Zhang et al. 2023) eliminate targeted unsafe concepts by guiding the denoising process away from unsafe concepts by fine-tuning the model. Even with carefully crafted prompts (Schramowski et al. 2023), these safety checkers and removal methods can prevent successful generation of NSFW images.
@@
Figure 1: An example on malicious prompt: sexy seductive little smile sophia vergara in nurse by agnes cecile enki bilal moebius, intricated details, lingerie, 3 / 4 back view, hair styled in a bun, bend over posture, full body portrait, extremely luminous bright design, pastel colours, drips, autumn lights. (a) Limitation of prior methods: Inconsistent semantics between NSFW generation and input prompt. (b) We precisely control the extent to which ‘nudity’ emerges in the generated images by a scalar $\lambda$.
@@
In the face of these safety measures, prior attack methods inevitably exhibit some drawbacks. As illustrated in Figure 1 (a), prior methods often compromise semantic fidelity to circumvent the robust safety checkers integrated into T2I generation models. Although they may succeed in generating NSFW images, the resulting content frequently fails to semantically align with the original input prompts. Furthermore, the challenge of searching an adversarial prompt within the discrete text space is exacerbated by the presence of safety checkers, leading to prolonged optimization processes or necessitating additional post-processing steps (Tsai et al. 2023; Zhuang, Zhang, and Liu 2023). These processes significantly extend the attack duration, thereby rendering such methods impractical.
@@
In this work, we show that the existing safety measures are not as effective as they may look. Our study is motivated by the lack of robustness in the embedding spaces (Zhuang, Zhang, and Liu 2023). We discover that the NSFW concepts are inherently embedded in the high-dimensional text embedding space. These NSFW concepts can be captured through specific concept embeddings, which allow us to manipulate the presence or absence of these concepts in generated images by simply adding or subtracting the corresponding embeddings. Furthermore, by adjusting the magnitude, we can precisely control the extent to which NSFW concepts emerge in the generated images, as shown in (b) of Figure 1. In light of these observations, we introduce Jailbreaking Prompt Attack (JPA), a more practical and universal attack in the black-box setting. JPA first searches a concept embedding that encapsulates the target malicious concept in the embedding space. These extra tokens are optimized to align with a group of antonyms of the target malicious concept generated by ChatGPT. To map back from the continuous embedding space to the discrete text token space, we optimize a prefix prompt concatenated at the beginning of the original prompt. To perform gradient ascent in discretized space, we propose a soft assignment with a gradient masking technique. The discrete token selection is modeled as a soft assignment, with the final token being chosen based on the highest probability. To prevent the appearance of sensitive words in the final prompt, we mask the gradients associated with a predefined list of NSFW terms. This ensures that only regular words achieve high probabilities during the soft assignment process.
@@
Bypass text-based checker (Jieli 2023; George 2023) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
Bypass image-based checker (Laion-ai 2023; Chhabra 2020) $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$
Bypass text-image based checker (Rando et al. 2022; CompVis 2022b) $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$
Bypass removal methods (Gandikota et al. 2023; Zhang et al. 2023) $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Applicable to online service (MidJourney 2023; Ramesh et al. 2022) $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Extensive experiments are performed with both open-sourced T2I models (CompVis 2022a) and closed-sourced online services (Ramesh et al. 2022; MidJourney 2023; Podell et al. 2023). Our JPA exhibits (1) controllable NSFW concept rendering (2) high semantic alignment with the target prompt (3) stealthiness to bypass all safety checkers. Equipped with our optimization technique, JPA also demonstrates a much faster speed than previous methods, executed in a fully automated manner.
Related Work
@@
T2I models with defense methods. To address concerns of NSFW generation, safety checkers is a common practice in existing T2I models. For example, Stable Diffusion (CompVis 2022a) filters out contents from 17 concepts (Rando et al. 2022) by leveraging cosine similarity between embeddings. DALL·E 2 (Ramesh et al. 2022) filters out contents from 11 categories such as hate, harassment, sexual, and self-harm. Midjourney (MidJourney 2023) blocks the generation of images that are not PG-13. Concept removal methods, on the other hand, force the model to forget the NSFW concepts in the generation. These methods have garnered widespread attention because when they can still generate images without NSFW content even if an unsafe text prompt in given. SLD (Schramowski et al. 2023) and SD-NP (Rombach et al. 2022) remove concepts during the inference stage. Other methods such as ESD (Gandikota et al. 2023) and FMN (Zhang et al. 2023) finetune the model to achieve concept removal. This ensures safety even in white-box setting, as the parameters have already been changed to remove the NSFW concepts.
@@
Adversarial attack on T2I models. Existing studies on adversarial attacks in Text-to-Image (T2I) models (Qu et al. 2023; Gao et al. 2023; Kou et al. 2023; Liu et al. 2023) have primarily focused on text modifications to identify functional vulnerabilities, without specifically targeting the generation of NSFW content. While some researchers (Rando et al. 2022) have explored techniques like prompt dilution to bypass safety checkers, these methods often depend on inherent patterns, limiting their portability and scalability. Efforts to assess the security of T2I models such as (Chin et al. 2024; Zhang et al. 2024) that induce offline models to produce NSFW content. However, these methods are ineffective for online services with unknown security checkers. QF-Attack (Zhuang, Zhang, and Liu 2023), Ring-A-Bell (Tsai et al. 2023), and SneakyPrompt (Yang et al. 2024) address this challenge by leveraging semantic information from the CLIP text encoder. Nevertheless, these approaches face two significant limitations: (1) they may fail to maintain consistent semantics with the target prompt and (2) often require lengthy search process to deliver the final adversarial prompt.
@@
Prompt perturbations in vision-language models. Due to the complexity of the text space, some studies employ prompt perturbation techniques to learn reusable prompts for generating specific images (Wen et al. 2023), or to investigate particular issues within T2I models, such as the hidden vocabulary in DALL·E 2 (Daras and Dimakis 2022; Maus et al. 2023). Other research uses fine-tuning methods to create prompts for generating personalized content (Gal et al. 2022). However, these approaches do not address the generation of NSFW content from a safety perspective.
Preliminary
@@
Defense Model
In this paper, we take into consideration all existing safety checkers, including text-based, image-based, and text-image-based checkers. Text-based safety checkers (Jieli 2023; George 2023) operate before image generation, detecting and filtering potentially unsafe words in the input text. Image-based safety checkers (Laion-ai 2023; Chhabra 2020) on the other hand, classify the generated images to determine whether they are unsafe, blocking those deemed inappropriate. Additionally, text-image-based safety checkers (Rando et al. 2022; CompVis 2022b) filter out NSFW content by calculating the cosine similarity between embeddings of malicious concepts and the image embedding under evaluation. This is usually achieved through a pre-defined threshold and images that goes above this threshold will be filtered. We also consider the concept removal methods (Schramowski et al. 2023; Gandikota et al. 2023; Zhang et al. 2023) eliminate targeted unsafe concepts by guiding the denoising process away from unsafe space by fine-tuning the model. We show later that these methods do not fully remove malicious concepts, which continue to reside within the high-dimensional embedding space.
Insights
@@
Following this setup, we then consider only prompts that (1) can map to unsafe images and (2) be sensitive to safety checkers. We denote a space that satisfies the above two criteria as $\mathcal{P}$ (red circle in the (b) of Figure 2), which can be divided into sensitive/insensitive space $\mathcal{P}=\mathcal{P}_{sensitive}\cup\mathcal{P}_{insensitive}$. The safety checker operates on the sensitive space $\mathcal{P}_{sensitive}$ and any prompt in it to either safe regions of image space or null space. For simplicity, we denote them all as $\mathcal{I}_{safe}$. Our goal is to bypass the safety checkers, i.e. find a mapping $J:\mathcal{P}_{sensitive}\rightarrow\mathcal{P}_{insensitive}$ and since $\mathcal{F}(\mathcal{P}_{insensitive})=\mathcal{I}_{unsafe}$, $J$ succeeds in attacking the T2I model. That is to say, for any $p_{t}\in\mathcal{P}_{sensitive}$, find a $p_{a}\in\mathcal{P}_{insensitive}$ that (1) goes crosses the decision boundary of the safety checkers and (2) semantically similar to $\mathcal{F}(p_{t})$. Then the research question becomes: how to find $J$ that can effectively and efficiently map $\mathcal{P}_{sensitive}$ to $\mathcal{P}_{insensitive}$? Our method stems from the conversation that high-dimensional text embedding space inherently contains NSFW concepts as concept embeddings. These embeddings can help us search a set of prefix tokens not in the sensitive space (typically looks like noise) that can be mapped to an NSFW image. Prefix optimization can help maintain the original semantics which fulfills the semantic fidelity requirement.
JPA: Jailbreaking Prompt Attack
@@
Figure 2: (a) Overview of the Jailbreaking Prompt Attack (JPA). Given a target prompt $p_{r}$ and a contrastive description of an NSFW concept <$r^{+},r^{-}$> such as <"nudity", "clothed"> for the "nudity" concept, we first obtain the embedding $\mathcal{T}(p_{r})$, which encapsulates both the semantic meaning and the unsafe concept. We then optimize our adversarial prompt $p_{a}$ to align $\mathcal{T}(p_{r})$ in the embedding space. (b) Equipped with safety checkers, the T2I model will map any prompt with sensitive words to either null space (w/ output) or a safe image (by concept removal). Our insight is to find an attacker function that map a sensitive prompt to an insensitive one while still maintaining NSFW content and semantic fidelity.
@@
In this section, we introduce our Jailbreaking Prompt Attack (JPA), as shown in Figure 2 (a). Given a sensitive prompt $p_{t}\in\mathcal{P}_{sensitive}$, JPA aims to search for an adversarial prompt $p_{a}\in\mathcal{P}_{insensitive}$. Denote a target prompt of length $n$ as $p_{t}=[p_{1},p_{2},\cdots,p_{n}]$, JPA starts by adding the $k$ learnable tokens at the beginning as the initial $p_{a}=[v_{1},...,v_{i},...,v_{k},p_{1},p_{2},\cdots,p_{n}]$. Each learnable token $v_{i}$ is first randomly selected from the vocabulary $V$, e.g. CLIP token vocabulary of 49,408 tokens. For each learnable token position $i$, any token from a vocabulary $V$ is considered a potential substitute. We enable the gradient of all $v_{i}$ and perform backpropagation on the attack learning objective to calculate the gradient.
@@
We optimize the $p_{a}$ with prefix learnable tokens to align with targeted NSFW concepts. Given a specific malicious concept, e.g. "nudity", we first generate $N$ pairs of antonyms associated with it utilizing ChatGPT4, e.g. "nude" and "clothed", denoted as $\{r^{+}\}^{N}$ and $\{r^{-}\}^{N}$. Then we subtract the antonyms embedding pairwisely and average them to capture the target NSFW concept in the embedding space.
@@
Then we add these embeddings to the original prompt embedding, denoted as $\mathcal{T}(p_{r})$. This effectively "adds" the NSFW concept to the original prompt.
@@
To search for $p_{a}$, we adopt the widely used PGD (Madry et al. 2017). However, since $p_{a}$ lies in the discrete text space, gradient accent is not directly applicable. We relaxed the strict one-hot selection into a soft assignment over the whole vocabulary. During the optimization, a softmax then weighted sum is used to generate embedding for each token.
@@
Despite its effectiveness and efficiency, this optimization process tends to generate highly sensitive words, as these words are closely aligned with NSFW concepts. To prevent detection by safety checkers, we propose masking the gradients of these sensitive words, ensuring they are not selected during optimization. Specifically, in each round of PGD, we assign extremely large gradient values (e.g., $91e$ 9) to these sensitive words. As a result, after the gradient updates, the probability values at these positions become very small, thereby reducing the likelihood of their selection. We detail our sensitive word list in Appendix. A detailed flow of JPA algorithm is presented in Appendix.
@@
Figure 3: Visualization results generated by JPA in NSFW concept under four online T2I services, (texts in red and black respectively represent the adversarial prompts from JPA and the original prompts from I2P datasets). We use Refer to caption and blur the displayed images for publication.
Experiment
Experimental setups.

### v4

@@ Abstract
We perform extensive experiments with open-sourced T2I models, e.g. (CompVis, 2022a) and closed-sourced online services, e.g.  (Ramesh et al., 2022; MidJourney, 2023; Podell et al., 2023) with black-box safety checkers. Results show that (1) JPA bypasses both text and image safety checkers (2) while preserving high semantic alignment with the target prompt. (3) JPA demonstrates a much faster speed than previous methods and can be executed in a fully automated manner. These merits render it a valuable tool for robustness evaluation in future text-to-image generation research. Our code is in https://github.com/mjc-ma/JPA
@@
Jailbreaking Prompt Attack:
A Controllable Adversarial Attack against Diffusion Models
Jiachen Ma1, Yijiang Li2, Zhiqing Xiao1, Anda Cao1, Jie Zhang3, Chao Ye1, Junbo Zhao1 1Zhejiang University,2UC San Diego, 3ETH Zurich mjc_zjdx@zju.edu.cn, j.zhao@zju.edu.cn
1 Introduction
@@
To mitigate these risks, contemporary T2I models incorporate stringent safety checkers and removal techniques to prevent the generation of NSFW content, including text-based (Jieli, 2023; George, 2023), image-based (Laion-ai, 2023; Chhabra, 2020) and (Rando et al., 2022; CompVis, 2022b) text-image-based safety checkers. These checkers either filter out sensitive words or classify unsafe images. Concept removal methods (Schramowski et al., 2023; Gandikota et al., 2023; Zhang et al., 2023) eliminate targeted unsafe concepts by guiding the denoising process away from unsafe concepts by fine-tuning the model. Even with carefully crafted prompts (Schramowski et al., 2023), these safety checkers and removal methods can prevent successful generation of NSFW images.
@@
Figure 1: An example on malicious prompt: “sexy seductive little smile sophia vergara in nurse by agnes cecile enki bilal moebius, intricated details, lingerie, 3 / 4 back view, hair styled in a bun, bend over posture, full body portrait, extremely luminous bright design, pastel colours, drips, autumn lights.” (a) Limitation of prior methods: Inconsistent semantics between NSFW generation and input prompt. (b) We precisely control the extent to which ‘nudity’ emerges in the generated images by a scalar $\lambda$.
@@
In the face of these safety measures, prior attack methods inevitably exhibit some drawbacks. As illustrated in Figure 1 (a), prior methods often compromise semantic fidelity to circumvent the robust safety checkers integrated into T2I generation models. Although they may succeed in generating NSFW images, the resulting content frequently fails to semantically align with the original input prompts. Furthermore, the challenge of searching an adversarial prompt within the discrete text space is exacerbated by the presence of safety checkers, leading to prolonged optimization processes or necessitating additional post-processing steps  (Tsai et al., 2023; Zhuang et al., 2023). These processes significantly extend the attack duration, thereby rendering such methods impractical.
@@
In this work, we show that the existing safety measures are not as effective as they may look. Our study is motivated by the lack of robustness in the embedding spaces (Zhuang et al., 2023). We discover that the NSFW concepts are inherently embedded in the high-dimensional text embedding space. These NSFW concepts can be captured through specific concept embeddings, which allow us to manipulate the presence or absence of these concepts in generated images by simply adding or subtracting the corresponding embeddings. Furthermore, by adjusting the magnitude, we can precisely control the extent to which NSFW concepts emerge in the generated images, as shown in (b) of Figure 1. In light of these observations, we introduce Jailbreaking Prompt Attack (JPA), a more practical and universal attack in the black-box setting. JPA first searches a concept embedding that encapsulates the target malicious concept in the embedding space. These extra tokens are optimized to align with a group of antonyms of the target malicious concept generated by ChatGPT. To map back from the continuous embedding space to the discrete text token space, we optimize a prefix prompt concatenated at the beginning of the original prompt. To perform gradient ascent in discretized space, we propose a soft assignment with a gradient masking technique. The discrete token selection is modeled as a soft assignment, with the final token being chosen based on the highest probability. To prevent the appearance of sensitive words in the final prompt, we mask the gradients associated with a predefined list of NSFW terms. This ensures that only regular words achieve high probabilities during the soft assignment process.
@@
Bypass text-based checker (Jieli, 2023; George, 2023) $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
Bypass image-based checker (Laion-ai, 2023; Chhabra, 2020) $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$
Bypass text-image based checker (Rando et al., 2022; CompVis, 2022b) $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$
Bypass removal methods (Gandikota et al., 2023; Zhang et al., 2023) $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Applicable to online service (MidJourney, 2023; Ramesh et al., 2022) $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Extensive experiments are performed with both open-sourced T2I models (CompVis, 2022a) and closed-sourced online services (Ramesh et al., 2022; MidJourney, 2023; Podell et al., 2023). Our JPA exhibits (1) controllable NSFW concept rendering, (2) high semantic alignment with the target prompt, and (3) stealthiness to bypass all safety checkers. Equipped with our optimization technique, JPA also demonstrates a much faster speed than previous methods, executed in a fully automated manner.
2 Related Work
T2I models with defense methods.
@@ 2 Related Work
To address concerns of NSFW generation, safety checkers is a common practice in existing T2I models. For example, Stable Diffusion (CompVis, 2022a) filters out contents from 17 concepts (Rando et al., 2022) by leveraging cosine similarity between embeddings. DALL·E 2 (Ramesh et al., 2022) filters out contents from 11 categories such as hate, harassment, sexual, and self-harm. Concept removal methods, on the other hand, force the model to forget the NSFW concepts in the generation. These methods have garnered widespread attention because when they can still generate images without NSFW content even if an unsafe text prompt in given. SLD (Schramowski et al., 2023) and SD-NP (Rombach et al., 2022) remove concepts during the inference stage. Other methods such as ESD (Gandikota et al., 2023) and FMN (Zhang et al., 2023) finetune the model to achieve concept removal. This ensures safety even in white-box setting, as the parameters have already been changed to remove the NSFW concepts.
Adversarial attack on T2I models.
@@
Existing studies on adversarial attacks in Text-to-Image (T2I) models (Qu et al., 2023; Gao et al., 2023; Kou et al., 2023; Liu et al., 2023) have primarily focused on text modifications to identify functional vulnerabilities, without specifically targeting the generation of NSFW content. While some researchers (Rando et al., 2022) have explored techniques like prompt dilution to bypass safety checkers, these methods often depend on inherent patterns, limiting their portability and scalability. Efforts to assess the security of T2I models such as (Chin et al., 2024; Zhang et al., 2024) that induce offline models to produce NSFW content. However, these methods are ineffective for online services with unknown security checkers. QF-Attack (Zhuang et al., 2023), Ring-A-Bell (Tsai et al., 2023), and SneakyPrompt (Yang et al., 2024) address this challenge by leveraging semantic information from the CLIP text encoder. Nevertheless, these approaches face two significant limitations: (1) they may fail to maintain consistent semantics with the target prompt and (2) often require lengthy search process to deliver the final adversarial prompt.
Prompt perturbations in vision-language models.
@@
Due to the complexity of the text space, some studies employ prompt perturbation techniques to learn reusable prompts for generating specific images (Wen et al., 2023), or to investigate particular issues within T2I models, such as the hidden vocabulary in DALL·E 2 (Daras and Dimakis, 2022; Maus et al., 2023). Other research uses fine-tuning methods to create prompts for generating personalized content (Gal et al., 2022). However, these approaches do not address the generation of NSFW content from a safety perspective.
3 Preliminary
@@
3.1 Defense Model
In this paper, we take into consideration all existing safety checkers, including text-based, image-based, and text-image-based checkers. Text-based safety checkers (Jieli, 2023; George, 2023) operate before image generation, detecting and filtering potentially unsafe words in the input text. Image-based safety checkers (Laion-ai, 2023; Chhabra, 2020) on the other hand, classify the generated images to determine whether they are unsafe, blocking those deemed inappropriate. Additionally, text-image-based safety checkers (Rando et al., 2022; CompVis, 2022b) filter out NSFW content by calculating the cosine similarity between embeddings of malicious concepts and the image embedding under evaluation. This is usually achieved through a pre-defined threshold and images that goes above this threshold will be filtered. We also consider the concept removal methods (Schramowski et al., 2023; Gandikota et al., 2023; Zhang et al., 2023) eliminate targeted unsafe concepts by guiding the denoising process away from unsafe space by fine-tuning the model. We show later that these methods do not fully remove malicious concepts, which continue to reside within the high-dimensional embedding space.
3.2 Insights
@@
Following this setup, we then consider only prompts that (1) can map to unsafe images and (2) be sensitive to safety checkers. We denote a space that satisfies the above two criteria as $\mathcal{P}$ (red circle in the Figure 2 (b)), which can be divided into sensitive/insensitive space $\mathcal{P}=\mathcal{P}_{\text{sensitive}}\cup\mathcal{P}_{\text{insensitive}}$. The safety checker operates on the sensitive space $\mathcal{P}_{\text{sensitive}}$ and any prompt in it to either safe regions of image space or null space. For simplicity, we denote them all as $\mathcal{I}_{\text{safe}}$. Our goal is to bypass the safety checkers, i.e. find a mapping $J:\mathcal{P}_{\text{sensitive}}\rightarrow\mathcal{P}_{\text{insensitive}}$ and since $\mathcal{F}(\mathcal{P}_{\text{insensitive}})=\mathcal{I}_{\text{unsafe}}$, $J$ succeeds in attacking the T2I model. That is to say, for any $p_{t}\in\mathcal{P}_{\text{sensitive}}$, find a $p_{a}\in\mathcal{P}_{\text{insensitive}}$ that (1) goes crosses the decision boundary of the safety checkers and (2) semantically similar to $\mathcal{F}(p_{t})$. Then the research question becomes: how to find $J$ that can effectively and efficiently map $\mathcal{P}_{\text{sensitive}}$ to $\mathcal{P}_{\text{insensitive}}$? Our method stems from the conversation that high-dimensional text embedding space inherently contains NSFW concepts as concept embeddings. These embeddings can help us search a set of prefix tokens not in the sensitive space (typically looks like noise) that can be mapped to an NSFW image. Prefix optimization can help maintain the original semantics which fulfills the semantic fidelity requirement.
4 JPA: Jailbreaking Prompt Attack
@@
Figure 2: (a) Overview of the Jailbreaking Prompt Attack (JPA). Given a target prompt $p_{r}$ and a contrastive description of an NSFW concept <$r^{+},r^{-}$> such as <“nudity”, “clothed”> for the “nudity” concept, we first obtain the embedding $\mathcal{T}(p_{r})$, which encapsulates both the semantic meaning and the unsafe concept. We then optimize our adversarial prompt $p_{a}$ to align $\mathcal{T}(p_{r})$ in the embedding space. (b) Equipped with safety checkers, the T2I model will map any prompt with sensitive words to either null space (w/ output) or a safe image (by concept removal). Our insight is to find an attacker function that map a sensitive prompt to an insensitive one while still maintaining NSFW content and semantic fidelity.
@@
In this section, we introduce our Jailbreaking Prompt Attack (JPA), as shown in Figure 2 (a). Given a sensitive prompt $p_{t}\in\mathcal{P}_{\text{sensitive}}$, JPA aims to search for an adversarial prompt $p_{a}\in\mathcal{P}_{\text{insensitive}}$. Denote a target prompt of length $n$ as $p_{t}=[p_{1},p_{2},\cdots,p_{n}]$, JPA starts by adding the $k$ learnable tokens at the beginning as the initial $p_{a}=[v_{1},...,v_{i},...,v_{k},p_{1},p_{2},\cdots,p_{n}]$. Each learnable token $v_{i}$ is first randomly selected from the vocabulary $V$, e.g. CLIP token vocabulary of 49,408 tokens. For each learnable token position $i$, any token from a vocabulary $V$ is considered a potential substitute. We enable the gradient of all $v_{i}$ and perform backpropagation on the attack learning objective to calculate the gradient.
@@
We optimize the $p_{a}$ with prefix learnable tokens to align with targeted NSFW concepts. Given a specific malicious concept, e.g. “nudity”, we first generate $N$ pairs of antonyms associated with it utilizing ChatGPT4, e.g. “nude” and “clothed”, denoted as $\{r^{+}\}^{N}$ and $\{r^{-}\}^{N}$. Then we subtract the antonyms embedding pairwisely and average them to capture the target NSFW concept in the embedding space.
@@
Then we add these embeddings to the original prompt embedding, denoted as $\mathcal{T}(p_{r})$. This effectively “adds” the NSFW concept to the original prompt.
@@
To search for $p_{a}$, we adopt the widely used PGD (Madry et al., 2017). However, since $p_{a}$ lies in the discrete text space, gradient accent is not directly applicable. We relaxed the strict one-hot selection into a soft assignment over the whole vocabulary. During the optimization, a softmax then weighted sum is used to generate embedding for each token.
@@
Despite its effectiveness and efficiency, this optimization process tends to generate highly sensitive words, as these words are closely aligned with NSFW concepts. To prevent detection by safety checkers, we propose masking the gradients of these sensitive words, ensuring they are not selected during optimization. Specifically, in each round of PGD, we assign extremely large gradient values (e.g., 1e9) to these sensitive words. As a result, after the gradient updates, the probability values at these positions become very small, thereby reducing the likelihood of their selection. We detail our sensitive word list in Appendix. A detailed flow of JPA algorithm is presented in Appendix.
@@
Figure 3: Visualization results generated by JPA in NSFW concept under four online T2I services, (texts in red and black respectively represent the adversarial prompts from JPA and the original prompts from I2P dataset).
5 Experiment
5.1 Experimental setups.
NSFW prompts.